The radii of 3 circles are 5cm, 7cm and 8cm as A as the biggest one and C as the smallest one and B in the middle. What is the perimeter of the triangle A to B to C
Answer by ikleyn(52786)

.
I guess that the circles are touching externally,

although the condition/your post does not state it.



Then one side of the triangle has the length of  5+7 = 12 cm;

the other side has the length of  5+8 = 13 cm

and the third side  has the length of  7+8 = 15 cm.


Then the perimeter of the triangle is  12 + 13 + 15 = 40 cm.


Answer.  The perimeter of the triangle ABC is 40 cm.